each of the numbers 0 through 9 is written on a piece of papaer and all of the pieces of paper are placed in a hat. one number is selected at random. determine the probability that the number selected is even
Answer by stanbon(75887)

each of the numbers 0 through 9 is written on a piece of papaer and all of the pieces of paper are placed in a hat. one number is selected at random. determine the probability that the number selected is even
---
Even numbers: 0,2,4,6,8,
Odd numbers: 1,3,5,7.9
---
P(even selected) = 5/10 = 1/2
=================================
cheers,
Stan H.
==================